OpenStudy (anonymous):

five times a number decreased by 18 minus 4 times the same number is -36. what is the number?

OpenStudy (anonymous):

5*x-18*4 = -36

OpenStudy (anonymous):

X= 7.2

OpenStudy (anonymous):

thank you so much SHAKEANDBAKE

OpenStudy (anonymous):

assume the number to be x now five times a number decreased by 18 means 5x -18 4 times the same number means 4x

OpenStudy (anonymous):

five times a number decreased by 18 minus 4 times the same number is -36.means 5x-18 -4x =-36

OpenStudy (anonymous):

YOU TOO matricked

OpenStudy (anonymous):

or 5x-4x = -36 +18 or x=-18

OpenStudy (anonymous):

LOL?